fifty heterozygous brown rabbits are placed with 100 homozygous white rabbits. what are the allele frequencies for brown fur (B) and for white fur (b)?


1
Expert's answer
2021-06-18T16:31:02-0400

Answer:

p = frequency of the B allele becomes 0.7

q = frequency of the b allele = 0.3

Explanation:

Hardy Weinberg equation is used to calculate frequencies.

Hardy Weinberg equation can be written as p² + 2pq + q² = 1

Where

p² = frequency of homozygous dominant

p = frequency of B dominant allele

q² = frequency of homozygous recessive

q = frequency of b recessive allele

2pq = frequency of heterozygous dominant

According to given data

p² = frequency of homozygous dominant = 0.33

p becomes √0.33

p = frequency of the B allele becomes 0.57

Sum of both frequencies = 100% = p + q = 1

so q = 1 - p

q = 1 - 0.57

q = frequency of the b allele = 0.43